ION IMAGING STUDIES OF CH_{2}I_{2} PHOTODISSOCIATION AT 248 NM

JULIA H. LEHMAN, \underline{HONGWEI LI} and MARSHA I. LESTER, Department of Chemsitry, University of Pennsylvania, Philadelphia, PA 19104-6323.

CH_{2}I_{2} plays an important role in atmospheric chemistry as a significant natural source of organohalide compounds. The photodissociation dynamics of CH_{2}I_{2} in the ultraviolet range of 277-305 nm via the two lowest B_{1} excited states has been well studied using one-color velocity map ion imaging (VMI) and photofragment translational spectroscopy. In this two-color experimental study, CH_{2}I_{2} is photodissociated by 248 nm via the B_{2} or A_{1} excited states to give rise to CH_{2}I and I (^{2}P_{3}_{/}_{2}) or I^{∗} (^{2}P_{1}_{/}_{2}). The iodine atoms are then state selectively ionized using a (2+1) resonance-enhanced multiphoton ionization process near 310 nm and detected by VMI. Preliminary results show about 85% of the available energy is being funneled into the internal energy of the CH_{2}I fragment, consistent with prior infrared emission results of Baughcum and Leone.^{a} The anisotropy parameter derived from the image indicates this is a fast dissociation process and reflects the character of the electronic transition. The internal energy distribution of the CH_{2}I fragment is of particular interest because of its subsequent reaction with O_{2} in a near thermo-neutral reaction to produce the smallest Criegee intermediate, CH_{2}OO. We anticipate that the internal energy contained in CH_{2}I will likely be carried into CH_{2}OO.

aS. L. Baughcum and S. R. Leone, J. Chem. Phys. 72, 6531 (1980).